A Complete Bibliography of Digital Government: Research and Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 August 2023 Version 1.06

Title word cross-reference

19 [ABK21, AAI21, BSA^{+}20, CN21, CGGV20, DMF^{+}21, DMS^{+}21, Fel21, Gai21, Gam21, GUJ21, GAB^{+}22, GWTC21, GBG^{+}20, GPCP20, HP21, JSH21, KLN22, KMJ20, KSM^{+}21, KGKS20, Kum22, LL21, Lun22, MP21, Red21, RAM^{+}21, Sie22, TRH21, Yaq21].

2020 [KKTT23].

Aarogya [GBG^{+}20]. Academic [DMS^{+}21]. Accessibility [VF22]. Accounting [WH20]. Addressing [HTH^{+}21]. Adoption [Fab22, TCMCAK23]. Africa [MF20]. African [KPR21]. Aftermath [TRH21].

Agencie [WH20]. Agents [KPR21]. Ahead [CN20, Sch20a]. Alternative [Sie22]. Americans [KPR21]. Amherst [DMS^{+}21]. Amid [LL21]. Analyses [YSP^{+}20]. Analysis [Bri21, Fab22, GBG^{+}20, LT22, MP21, PLP21, SAVC20, Smy20, SGC22, Yaq21]. Analyzing [GJ22, HEMJ20]. Annotation [RKG20]. Anti [TH22]. Anti-Corrosion [TH22]. Application [GBG^{+}20, SGC22]. Approach [ABK21]. ARIMA [KGKS20]. Artificial [Lee23, SM21]. Assembly [GAB^{+}22]. Assistance [AGH^{+}21]. assisted [RKG20]. Austria [PLP21]. Authoritarian [Sch22]. Automated [HTH^{+}21, RBDB21]. Automatic [BSA^{+}20]. Avoid [Mul22].

Back [CN20, HMW^{+}20, Sch20a]. Based [Bri21, Fel21, PXW22, TH22, Guo22, LT22, YSP^{+}20]. Behavior [LL21]. Being [Bez20]. Beth [Kam22, Scl20]. Better [ACVP^{+}21, Mul22]. between [MF20, SLO20]. Beyond [GWTC21]. Big [LT22]. Blockchain [Bri21]. Blockchain-Based [Bri21]. Blockchains [CDZ^{+}20]. Book [Kam22]. Both [Mul22]. Bots [Per20]. Brain [Mul22]. Branch [FM22]. Bridging [KPR21]. Brings [JSH21]. Building [Gil22, LM22, SPK^{+}22]. Burden [HJY^{+}21, Smy21]. Business [TRH21].

C [KA23]. C-suite [KA23]. Campaign [SAVC20]. Can [AGH^{+}21, Lee23, Mul22, Sch20b]. Canvas [GJ22]. Canvas-Analyzing [GJ22]. Case [CGGV20, GAB^{+}22, GBG^{+}20, Gup21, HJY^{+}21, HTH^{+}21]. Cases [CDZ^{+}20, Lee23]. Cast [SGC22]. Centre [RAM^{+}21]. Centuries [Bra22]. Challenges [CDZ^{+}20, Gup21, JSH21, SPK^{+}22]. Change [ALH20, Kam22]. Chief [NC20]. Children [GWTC21]. China [LL22, LM22]. Citizen [ACVP^{+}21, GAB^{+}22, HEMJ20, LL22, RBDB21]. Citizens [CSFV22, CGGV20, SM21, TCMCAK23]. City [ASR21, RAM^{+}21]. Civic [Gil22, LL22, RDM21]. Civil [GD20]. Cloud [AGH^{+}21]. Co [PYM22]. Co-tailoring [PYM22]. Coalition [SLO20]. Code [JS20]. Collaborative [CGGV20, GT22, Per20]. Collective [Lun22, MN22, Mul22, SPK^{+}22]. Colombia [DDRF21]. Commentary [CN21, LL21]. Communication [KLN22]. Communications [HNSK20]. Communities [PMM^{+}22]. Community [DMS^{+}21]. Compared [CSFV22]. Computing [Sch22]. Conceptualizing [ASR21]. Concerns [Red21]. conducted [DMS^{+}21]. Conferences [PMM^{+}22]. Confirmation [Guo22]. Conflicting [MF20]. Congress [KLN22]. Connecting [KLN22]. Consortium [PYM22]. Constitutions [Sie22]. Consumption [GT22]. Contact [BMS22]. Contests [Gam21]. Convening [GAB^{+}22]. Conversational [KPR21]. Conversations [WH20]. Cope [Gai21]. Corporation [GD20]. Corrosion [TH22]. Costs [HH22]. Could [GD20, PYM22]. Country [HMW^{+}20]. County [MP21]. Court [HJY^{+}21]. CoV2 [Smy21]. COVID [RBDB21, ABK21, AAI21, BSA^{+}20, CN21, CGGV20, DMF^{+}21, DMS^{+}21, Fel21, Gam21, GUJ21, GAB^{+}22, GWTC21, GBG^{+}20, GPCP20, HP21, JSH21, KLN22, KMJ20, KSM^{+}21, KGKS20, Kum22, LL21, Lun22, MP21, Red21, RAM^{+}21, Sie22, TRH21, Gai21, Yaq21]. COVID-19 [RBDB21, ABK21, AAI21, BSA^{+}20, CN21, CGGV20, DMF^{+}21, DMS^{+}21, Fel21, Gam21, GUJ21, GAB^{+}22, GWTC21, GBG^{+}20, GPCP20, HP21, JSH21, KLN22, KMJ20, KSM^{+}21, KGKS20, Kum22, LL21, Lun22, MP21, Red21, RAM^{+}21, Sie22, TRH21, Gai21, Yaq21]. Create [Lee23, Mul22]. Creation [GJ22]. Crisis [AGH^{+}21, CGGV20, Gil22, HNSK20, KLN22, Lun22, Mar22, Sie22]. Critical [Kum22]. Crowdlaw [MN22, Sie22]. Crowdsource [Gam21]. Crowdsourcing [FM22, Mar22, RDM21]. CT [Kam22]. Cyber [RAM^{+}21, TRH21]. Cyber-physical [RAM^{+}21]. Cyber-Security [TRH21].

Dashboard [RBDB21]. Dashboards

[CSFV22]. Data [Bez20, BSA^{+}20, CSFV22, DDRF21, GJ22, HP21, JSH21, LT22, Lun22, RBDB21, Smy20, YSP^{+}20]. Data-Driven [HP21]. Dealing [TRH21]. Decision [HP21]. Decisions [HTH^{+}21]. Deconstructing [RKG20]. Deep [ABK21]. Delivering [AGH^{+}21]. Democracy [ACVP^{+}21, CN20, GD20, NC20, Per20, Sch20b, Scl20]. Democratic [Mar22, Per20].

Demographic [MP21]. Department [HH22]. Design [Bez20, GT22, RDM21, Sgo22]. Detailed [Gai21]. Detection [BSA^{+}20]. Developing [GMP20]. Development [GMP20, Gup21, KS21]. Dialogic [WH20]. Differences [SLO20]. Digital [ALH20, CN20, GD20, GT22, Grö20, Guo22, HMW^{+}20, KA23, LL22, LM22, NC20, Per20, Sch20a]. Discourse [HP21]. Disease [GMS^{+}21]. Disparity [KPR21]. Distancing [JYAM^{+}20]. Divertible [HH22]. Domain [Sch20a]. Driven [HP21]. During [GAB^{+}22, GWTC21, HNSK20, KLN22, KSM^{+}21, RBDB21, Yaq21, CGGV20, KMJ20, Kum22].

E-governance [GMP20]. E-government [Mku23, TCMCAK23]. e-Rulemaking [Per20]. e-Service [PYM22]. Economic [LT22]. Economy [MF20]. Editorial [CN20, CN21]. Education [PXW22]. Efforts [Lun22]. Election [SGC22, YSP^{+}20]. Emergencies [Mul22]. Emergency [HNSK20, HH22, MN22]. Emerging [KKTT23]. Emotional [GWTC21]. Employees [ALH20]. Empower [FM22]. Engage [CGGV20]. Engagement [GD20, LL22]. Engagement-Evidence [LL22]. Engaging [CSFV22]. Engineering [Gup21]. Episode [GD20]. Equity [JSH21]. Essential [AGH^{+}21]. Estimating [Smy21]. Ethical [Kum22]. Europe [Bri21]. Evaluation [PXW22, VF22]. Evangelist [NC20]. Events [DDRF21]. Evidence [LL22]. Examining [WH20]. Expectation [Guo22]. Expertise [FM22]. Explanations [HTH^{+}21]. Explorative [Bri21]. Exploring [Lun22, Mku23].

Facebook [SAVC20, SLO20, WH20]. facilitated [Red21]. Facilitating [PMM^{+}22]. Factors [TCMCAK23]. Fair [Bez20, Bez20]. Fascinating [Sch20a].

Fatalities [MP21]. Fatality [Smy21]. Fear [DMF^{+}21]. Federal [JS20]. Fighting [GUJ21]. Finding [VF22]. Fix [Kam22]. Fore [JSH21]. Framework [JS20, RDM21]. Future [Lun22, NC20, Per20]. Fuzzy [PXW22].

G [NC20]. Gaming [KMJ20]. Gathering [RBDB21]. Getting [HMW^{+}20]. Ghana [MF20]. Global [KS21, PMM^{+}22, RBDB21, SPK^{+}22]. go [PYM22]. Google [NC20]. Governance [Bri21, Kum22, GMP20]. Government [ALH20, AGH^{+}21, ASR21, CSFV22, CN21, CDZ^{+}20, GD20, GT22, GPCP20, HEMJ20, Kam22, KA23, LM22, LT22, Mul22, PXW22, SAUK20, Sch20a, VF22, Mku23, TCMCAK23]. Government-Citizen [HEMJ20]. Governmental [PYM22]. Governments [CGGV20, Mul22]. Guide [Kam22].

Hackathons [Gam21]. Happens [Smy20]. Haven [Kam22]. Health [HNSK20, HP21, KPR21]. High [TCMCAK23]. High-level [TCMCAK23]. Hindcasting [DDRF21]. Hope [GD20]. Horizon [KKTT23]. Hours [Bra22]. Human [RKG20]. Human-assisted [RKG20].

ICCC [RAM^{+}21]. Identities [GUJ21]. IDIOMS [GMS^{+}21]. Images [ABK21]. Imaging [GMS^{+}21]. Impact [DMF^{+}21, DMS^{+}21, GUJ21]. Impacts [CN21, LL22]. Implementations [Bri21]. Implementing [Sgo22]. Implications [Fel21]. Improvement [TH22]. Index [GMP20]. India [GBG^{+}20, KGKS20, Kum22]. Individual [CSFV22]. Infectious [GMS^{+}21]. Influential [TCMCAK23]. Information [Fel21, Gai21, HJY^{+}21, Lun22]. Information-Based [Fel21].

Infrastructure [RAM^{+}21]. Innovation [Gam21, Mku23]. Innovative [PXW22]. Insights [Gup21]. Institutional [Fel21]. Integrated [Guo22]. Integrating [KMJ20, RDM21]. Intelligence [Lee23, Lun22, MN22, SPK^{+}22]. Interactions [HEMJ20, KMJ20]. Internal [PLP21]. Internet [NC20, DDRF21]. Interpersonal [Fel21]. Interrelations [ASR21]. Interview [NC20, Scl20]. Introduction [LM22, MN22, SAUK20]. Invisible [TRH21]. Iraq [AAI21]. Issue [LM22, MN22, SAUK20]. Issues [CN21]. Italy [KGKS20, VF22]. Itself [Jar20]. IV [GD20].

Japan [KGKS20]. Journey [Bra22].

Keep [AGH^{+}21]. Knowledge [JS20].

Leaders [SM21]. Leadership [KA23]. Learned [KMJ20, Lee23, PMM^{+}22]. Learning [ABK21, ACVP^{+}21, GWTC21, GPCP20]. Legislative [FM22, RKG20]. Legitimacy [Mar22]. Lessons [KMJ20, PMM^{+}22, Lee23]. Level [MP21, TCMCAK23]. like [RAM^{+}21]. Literacy [LL22]. Literature [TCMCAK23]. Local [VF22]. Location [BMS22, YSP^{+}20]. Location-based [YSP^{+}20]. Lockdowns [Smy20]. Look [CN20]. Looking [Sch20a]. LOP [TH22]. LSTM [KGKS20].

Machine [ACVP^{+}21, GPCP20]. Making [HP21, HEMJ20]. Management [Mar22]. Massachusetts [DMS^{+}21]. Mathematics [PXW22]. Meaningful [KMJ20]. Media [CGGV20, HNSK20, SAUK20]. Mediation [HJY^{+}21]. Medicaid [HH22]. Medical [KS21]. Meetings [KMJ20]. Members [SLO20]. Messages [Yaq21]. Meta [Sgo22]. Meta-simulation [Sgo22]. Method [LT22, PXW22]. Methodology [Sgo22].

Migration [PLP21]. Minipublic [GAB^{+}22]. Mobile [GBG^{+}20]. Mobility [JYAM^{+}20]. Model [Guo22, KGKS20, TCMCAK23]. Modeling [KSM^{+}21]. Models [Gam21, GPCP20]. Monitoring [GMS^{+}21, RAM^{+}21]. MPs [SLO20]. Multiple [Gup21]. Municipal [Fab22]. Municipalities [PYM22].

National [Gai21]. Nations [GMP20]. Need [AGH^{+}21]. Nerve [RAM^{+}21]. Network [PLP21, SGC22]. Next [Smy20]. Normal [Mul22]. Noveck [Kam22, Scl20]. Novel [Sgo22].

Occupancy [JYAM^{+}20]. Online [LL21]. Open [CSFV22, MMBM21, GJ22, Lun22, Mku23, RDM21]. Opportunities [SPK^{+}22]. Opposition [SLO20]. Oppression [Sch22]. Oregon [GAB^{+}22]. Organizations [GT22]. Our [Kam22]. Outbreak [GMS^{+}21, GPCP20]. Oversight [FM22].

Pandemic [DMS^{+}21, Gai21, GUJ21, GAB^{+}22, GWTC21, KMJ20, KSM^{+}21, Kum22, KS21, LL21, MF20, RBDB21, Yaq21]. Pandemics [Lun22, RAM^{+}21]. Participating [Grö20]. Participation [ACVP^{+}21, RDM21]. Passion [DMF^{+}21]. Past [NC20]. Pattern [MP21]. Penetration [BMS22]. Performance [BMS22]. Perspective [Kum22]. pgs [Kam22]. Phase [BSA^{+}20]. Physical [PXW22, RAM^{+}21]. Pilot [GAB^{+}22]. Platform [GT22, PYM22]. Platforms [Lun22]. Play [MMBM21]. Pneumonia [ABK21]. Police [WH20]. Policies [CN21, Gai21, Kum22]. Policy [HEMJ20, JYAM^{+}20, Sgo22]. Policy-making [HEMJ20]. Policymaking [FM22]. Political [KLN22, SAVC20]. Politics [HP21]. Politika [Sgo22]. Poses [Sch22]. Position [BMS22]. Posts [SAVC20]. Potential [GWTC21]. Power [Gil22]. Practical [Kam22]. Practice [Bra22, Sch20a]. Practices [Mku23, WH20]. Predict [GPCP20, MP21]. Predicting [HH22, HJY^{+}21]. Prediction [KGKS20, LT22]. Preface [CN21]. Preparedness [Lun22]. Present [NC20]. President [NC20, Yaq21]. Press [Kam22]. Privacy [BMS22]. Problems [Kam22]. Proceedings [RKG20]. Processes [MMBM21]. Product [KS21]. Products [KS21]. Programs [AGH^{+}21]. Projecting [AAI21]. Projects [KKTT23]. Properties [HJY^{+}21]. Protection [TH22]. Provenance [HTH^{+}21]. Providing [Gai21]. Provision [RBDB21]. Psychological [DMF^{+}21]. Psychosocial [DMF^{+}21]. Public [Bri21, MMBM21, GD20, HP21, HEMJ20, KKTT23, Kam22, Lee23, RDM21, Sgo22]. Pulse [KSM^{+}21].

Quality [PXW22].

Rapid [CN21, Gam21, KS21]. ray [ABK21]. Reacting [SAVC20]. Real [SM21]. Realities [Bri21]. Rebounds [Smy20]. Reciprocities [ASR21]. Reconstruction [Guo22]. Recovery [GAB^{+}22, Sie22]. Reducing [HJY^{+}21]. Reflections [PMM^{+}22]. Regional [GMP20]. Regulations [JS20]. Regulatory [HTH^{+}21]. Relearning [Jar20]. Remember [Mul22]. Representation [JS20, KLN22]. Requests [HJY^{+}21]. Requirement [Gup21]. Requirements [HTH^{+}21]. Research [Bra22, Sch20a]. Response [Lun22, Red21]. Responses [CN21, Gam21]. Responsible [RBDB21]. Results [DMS^{+}21]. Review [KKTT23, Kam22, TCMCAK23]. Rich [JS20]. Richard [Scl20]. Risk [HNSK20]. Risks [TRH21]. Role [GWTC21, Lun22]. Rulemaking [Per20]. Running [AGH^{+}21].

Safely [HMW^{+}20]. Safety [TH22]. SARS [Smy21]. SARS-CoV2 [Smy21]. Scalable [AGH^{+}21]. Science [FM22, GT22]. Sclerosis [ALH20]. Sclove [Scl20]. Screening [ABK21]. Sector [Gup21, KKTT23]. Security [TRH21]. Self [GUJ21]. Self-sovereign [GUJ21]. Semantically [JS20]. Sentiment [SAVC20, YSP^{+}20, Yaq21]. Service [PYM22, TCMCAK23]. Setu [GBG^{+}20]. Sharing [Lun22]. Short [AAI21]. Short-Term [AAI21]. Signals [KSM^{+}21]. Simone [Kam22, Scl20]. simulation [Sgo22]. Situation [LT22]. Small [TRH21]. Smart [ASR21, RAM^{+}21]. Smartphone [SGC22]. Smokestack [TH22]. Social [CGGV20, GWTC21, Gup21, HNSK20, JYAM^{+}20, KSM^{+}21, SAUK20]. Social-Emotional [GWTC21]. Societies [SPK^{+}22]. Society [GD20, Grö20, Guo22, Jar20]. Socio [SGC22]. Socio-technical [SGC22]. Software [Gup21, GD20]. Solution [HMW^{+}20]. Solutions [AGH^{+}21]. Solving [Kam22]. sovereign [GUJ21]. Spain [CGGV20, KGKS20]. Special [CN21, LM22, MN22, SAUK20]. Spreading [KGKS20]. Standards [Bez20]. Strategies [MF20, TH22]. Studies [Gup21, TH22]. Study [GAB^{+}22, GBG^{+}20, HTH^{+}21]. Stupidity [Mul22]. Success [HJY^{+}21]. Successful [Gam21]. suite [KA23]. Support [HEMJ20, Sch20b]. Supporting [AGH^{+}21]. survey [DMS^{+}21]. Survival [Fab22]. System [GMS^{+}21].

tailoring [PYM22]. Tale [Fab22]. Talk [Jar20]. Tanzania [Mku23, MF20]. technical [SGC22]. Technologies [CGGV20, Fab22, KKTT23]. Technology [FM22, GWTC21, Red21, Sch20b, Sch22, Scl20]. Technology-facilitated [Red21]. Teleworking [KMJ20]. Term [AAI21]. Textual [HJY^{+}21]. Theory [Guo22].

Things [Lee23]. Think [Mul22]. Thinking [RDM21]. Those [AGH^{+}21]. Threat [Sch22]. Three [Fab22]. Timely [RBDB21]. Times [AGH^{+}21, Mul22]. Tool [GBG^{+}20]. Tracing [BMS22]. Tracking [GBG^{+}20, JYAM^{+}20, RAM^{+}21]. Tradeoffs [Red21]. Training [PXW22]. Transcription [RKG20]. Transform [GD20]. Traps [Mul22]. Trend [AAI21]. Trends [LM22]. Trump [Yaq21]. Trust [Fel21, Guo22]. Turn [SM21]. Tweeting [Yaq21]. Twitter [Fab22, HEMJ20, YSP^{+}20, Yaq21]. Two [KLN22, PMM^{+}22]. Two-Way [KLN22].

UK [KGKS20]. Understand [Mul22]. Understanding [KKTT23]. Unemployment [AGH^{+}21]. University [Kam22, DMS^{+}21]. Unraveling [KSM^{+}21]. Urban [JYAM^{+}20, TH22]. Use [CDZ^{+}20, HNSK20, KKTT23, Lee23]. User [Red21]. Users [SAVC20]. Using [CGGV20, DDRF21, HJY^{+}21, KGKS20, MP21, GT22]. Utilizing [RAM^{+}21].

Value [CSFV22, GJ22, Lee23]. Vice [NC20]. Video [RKG20]. Vinton [NC20]. Violent [DDRF21]. Viral [ABK21]. Virtual [KMJ20, PMM^{+}22]. Visualization [BSA^{+}20, YSP^{+}20]. Visualizations [CSFV22]. Visualizing [HEMJ20]. Vote [SGC22]. vs [Sie22].

Way [KLN22]. Web [Sgo22]. Websites [Fab22, VF22]. Wind [ALH20]. Work [HMW^{+}20]. Workplace [Fel21]. World [Kam22]. Worldwide [Sch22]. Would [SM21]. wrong [PYM22].

X [ABK21]. X-ray [ABK21]. Xenophobic [LL21].

Yale [Kam22]. Years [CN20]. Young [GWTC21]. YouTube [Fab22].

Zika [HNSK20].

References

Aldeer:2021:PST

[AAI21] Murtadha Aldeer, Ahmed Al Hilli, and Issam S. Ismail. Projecting the short-term trend of COVID-19 in Iraq. Digital Government: Research and Practice (DGOV), 2(1): 14:1–14:7, January 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                                     URL https://dl.acm.org/
                                                     doi/10.1145/3431769.

Ahmed:2021:DLA

[ABK21] Faizan Ahmed, Syed Ahmad Chan Bukhari, and Fazel Keshtkar. A deep learning approach for COVID-19 8 viral pneumonia screening with X-ray images. Digital Government: Research and Practice (DGOV), 2(2):18:1–18:12, March 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                                     //dl.acm.org/doi/10.1145/
                                                     3431804.

Arana-Catania:2021:CPM

[ACVP^{+}21] Miguel Arana-Catania, Felix-Anselm Van Lier, Rob Procter, Nataliya Tkachenko, Yulan He, Arkaitz Zubiaga, and Maria Liakata. Citizen participation and machine learning for a better democracy. Digital Government: Research and Practice (DGOV), 2(3):27:1– 27:22, July 2021. CODEN ????

ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3452118.

Angell:2021:DUA

[AGH^{+}21] Mintaka Angell, Samantha Gold, Mark Howison, Victoria Kidd, Daniel Molitor, Casey Burns, Chris Johnson, Matthew Kahn, Stuart Venzke, Sandra Deneault, Donald Doweiko, Stephen Dziembowski, Bijay Kumar, Patrick O’donnell, Chirag Patel, Andrew Reidl, Brian Tardiff, Justine S. Hastings, Scott Jensen, Angelika Pellegrino, Amelia Roberts, and Rahul Sarathy. Delivering unemployment assistance in times of crisis: Scalable cloud solutions can keep essential government programs running and supporting those in need. Digital Government: Research and Practice (DGOV), 2(1):5:1–5:11, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428125.

Andersen:2020:DSW

[ALH20] Kim Normann Andersen, Jungwoo Lee, and Helle Zinner Henriksen. Digital sclerosis? Wind of change for government and the employees. Digital Government: Research and Practice (DGOV), 1(1): 9:1–9:14, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3360000.

Anthopoulos:2021:CSG

[ASR21] Leonidas Anthopoulos, Kleanthis Sirakoulis, and Christopher G. Reddick. Conceptualizing smart government: Interrelations and reciprocities with smart city. Digital Government: Research and Practice (DGOV), 2(4):33:1–33:28, October 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3465061.

Bezuidenhout:2020:BFA

[Bez20] Louise Bezuidenhout. Being fair about the design of FAIR data standards. Digital Government: Research and Practice (DGOV), 1(3):18:1–18:7, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3399632.

Bedogni:2022:LCT

[BMS22] Luca Bedogni, Federico Montori, and Flora Salim. Location contact tracing: Penetration, privacy, position, and performance. Digital Government: Research and Practice (DGOV), 3(3):18:1–18:??, July 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532669.

Bragge:2022:CHJ

[Bra22] Peter Bragge. From centuries to hours: The journey of research into practice. Digital Government: Research and Practice (DGOV), 3(2): 16:1–16:??, April 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3529166.

Brinkmann:2021:RBB

[Bri21] Maik Brinkmann. The realities of blockchain-based new public governance: an explorative analysis of blockchain implementations in Europe. Digital Government: Research and Practice (DGOV), 2(3):29:1– 29:14, July 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462332.

Biswas:2020:CDV

[BSA^{+}20] Pradipta Biswas, Kamalpreet Singh Saluja, Somnath Arjun, Lrd Murthy, Gowdham Prabhakar, Vinay Krishna Sharma, and Jeevitha Shree Dv. COVID-19 data visualization through automatic phase detection. Digital Government: Research and Practice (DGOV), 1(4):25:1–25:8, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3411756.

Clavin:2020:BGU

[CDZ^{+}20] James Clavin, Sisi Duan, Haibin Zhang, Vandana P. Janeja, Karuna P. Joshi, Yelena Yesha, Lucy C. Erickson, and Justin D. Li. Blockchains for government: Use cases and challenges. Digital Government: Research and Practice (DGOV), 1(3):22:1–22:21, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3427097.

Criado:2020:UCT

[CGGV20] J. Ignacio Criado, Ariana Guevara-Gómez, and Julián Villodre. Using collaborative technologies and social media to engage citizens and governments during the COVID-19 crisis. The case of Spain. Digital Government: Research and Practice (DGOV), 1(4):30:1– 30:7, December 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3416089.

Chun:2020:EDD

[CN20] Soon Ae Chun and Beth Simone Noveck. Editorial: Digital democracy: 25 years look back, look ahead. Digital Government: Research and Practice (DGOV), 1(1):1:1– 1:3, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3374904.

Chun:2021:EPS

[CN21] Soon Ae Chun and Beth Simone Noveck. Editorial preface special issues on COVID-19 commentary: Rapid responses, government policies, and impacts. Digital Government: Research and Practice (DGOV), 2(3):25e:1– 25e:4, July 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3463837.

Chokki:2022:ECO

[CSFV22] Abiola Paterne Chokki, Anthony Simonofski, Benoît Frénay, and Benoît Vanderose. Engaging citizens with open government data: The value of dashboards compared to individual visualizations. Digital Government: Research and Practice (DGOV), 3(3):21:1– 21:??, July 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558099.

Daughton:2021:HVE

[DDRF21] Ashlynn R. Daughton, Sara Y.

Del Valle, Chrysm Watson Ross, and Geoffrey Fairchild. Hindcasting violent events in Colombia using Internet data. Digital Government: Research and Practice (DGOV), 2(3): 28:1–28:14, July 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3462211.

Dey:2021:CPP

[DMF^{+}21] Nilanjan Dey, Rishabh Mishra, Simon James Fong, K. C. Santosh, Stanna Tan, and Rubén González Crespo. COVID-19: Psychological and psychosocial impact, fear, and passion. Digital Government: Research and Practice (DGOV), 2(1):3:1–3:4, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428088.

Deznabi:2021:ICP

[DMS^{+}21] Iman Deznabi, Tamanna Motahar, Ali Sarvghad, Madalina Fiterau, and Narges Mahyar. Impact of the COVID-19 pandemic on the academic community results from a survey conducted at University of Massachusetts Amherst. Digital Government: Research and Practice (DGOV), 2(2): 22:1–22:12, March 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3436731.

Faber:2022:TTT

[Fab22] Bram Faber. A tale of three technologies: a survival analysis of municipal adoption of Websites, Twitter, and YouTube. Digital Government: Research and Practice (DGOV), 3(3):20:1–20:??, July 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3559108.

Fell:2021:TCI

[Fel21]  Lauren Fell.  Trust and

COVID-19: Implications for interpersonal, workplace, institutional, and informationbased trust. Digital Government: Research and Practice (DGOV), 2(1):10:1–10:5, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428472.

Fisher:2022:CST

[FM22] Michael A. Fisher and Lindsay K. Milliken. Crowdsourcing science and technology expertise to empower legislative branch oversight and policymaking. Digital Government: Research and Practice (DGOV), 3(2):11:1–11:??, April 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3524065.

Gastil:2022:CMD

[GAB^{+}22] John Gastil, Chris Anderson, Laura Black, Stephanie Burkhalter, Soo-Hye Han, Justin Reedy, Robert Richards, and John Rountree. Convening a minipublic during a pandemic: a case study of the Oregon citizens’ assembly pilot on COVID-19 recovery. Digital Government: Research and Practice (DGOV), 3(2): 13:1–13:??, April 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3524064.

Gaie:2021:PDI

[Gai21] Christophe Gaie. Providing detailed information on national policies to cope with the Covid-19 pandemic. Digital Government: Research and Practice (DGOV), 2(1): 7:1–7:11, January 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428089.

Gama:2021:SMH

[Gam21] Kiev Gama. Successful models of hackathons and innovation contests to crowdsource rapid responses to COVID-19. Digital Government: Research and Practice (DGOV), 2(2): 20:1–20:7, March 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3431806.

Gupta:2020:ACT

[GBG^{+}20] Rajan Gupta, Manan Bedi, Prashi Goyal, Srishti Wadhera, and Vaishnavi Verma. Analysis of COVID-19 tracking tool in India: Case study of aarogya setu mobile application. Digital Government: Research and Practice (DGOV), 1(4):28:1– 28:8, December 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3416088.

Gastil:2020:DDE

[GD20] John Gastil and Todd Davies.

Digital democracy: Episode IV — a new hope: How a Corporation for Public Software could transform digital engagement for government and civil society. Digital Government: Research and Practice (DGOV), 1(1):6:1–6:15, February 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3342194.

Gilman:2022:BCP

[Gil22] Hollie Russon Gilman. Building civic power in crisis. Digital Government: Research and Practice (DGOV), 3(2): 10:1–10:??, April 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3529164.

Gao:2022:ODC

[GJ22] Yingying Gao and Marijn Janssen. The open data canvas-analyzing value creation from open data. Digital Government: Research and Practice (DGOV), 3(1): 5:1–5:15, January 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3511102.

Gupta:2020:RGD

[GMP20] Rajan Gupta, Sunil Kumar Muttoo, and Saibal Kumar Pal. Regional e-governance development index for developing nations. Digital Government: Research and Practice (DGOV), 1(3):20:1–20:26, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3386163.

Gangopadhyay:2021:IID

[GMS^{+}21] Aryya Gangopadhyay, Michael Morris, Babak Saboury, Eliot Siegel, and Yelena Yesha. ID-IOMS: Infectious disease imaging outbreak monitoring system. Digital Government: Research and Practice (DGOV), 2(1):15:1–15:5, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428092.

Gupta:2020:MLM

[GPCP20] Rajan Gupta, Gaurav Pandey, Poonam Chaudhary, and Saibal K. Pal. Machine learning models for government to predict COVID-19 outbreak. Digital Government: Research and Practice (DGOV), 1(4): 26:1–26:6, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3411761.

Gronlund:2020:PDS

[Grö20] Åke Grönlund. Participating in the digital society. Digital Government: Research and Practice (DGOV), 1(2): 17:1–17:13, April 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3361865.

Gebeyehu:2022:CCD

[GT22] Seffi Gebeyehu and Hossana Twinomurinzi. A collaborative consumption digital platform for government organizations using design science. Digital Government: Research and Practice (DGOV), 3(1): 4:1–4:14, January 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3510615.

Gans:2021:SSI

[GUJ21] Rachel Benchaya Gans, Jolien Ubacht, and Marijn Janssen. Self-sovereign identities for fighting the impact of COVID-19 pandemic. Digital Government: Research and Practice (DGOV), 2(2):16:1–16:4, March 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429629.

Guo:2022:DTR

[Guo22] Yuanyuan Guo. Digital trust and the reconstruction of trust in the digital society: an integrated model based on trust theory and expectation confirmation theory. Digital Government: Research and Practice (DGOV), 3(4):26:1–26:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3543860.

Gupta:2021:REC

[Gup21] Varun Gupta. Requirement engineering challenges for social sector software development: Insights from multiple case studies. Digital Government: Research and Practice (DGOV), 2(4):34:1–34:13, October 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3479982.

Goagoses:2021:PRT

[GWTC21] Naska Goagoses, Heike Winschiers-Theophilus, and Michael Bosomefi Chamunorwa. The (potential) role of technology for young children’s socialemotional learning: During and beyond the COVID-19 pandemic. Digital Government: Research and Practice (DGOV), 2(1):13:1–13:6, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429986.

Hubert:2020:AVG

[HEMJ20] Rocío B. Hubert, Elsa Estevez, Ana Maguitman, and Tomasz Janowski. Analyzing and visualizing government-citizen interactions on Twitter to support public policy-making. Digital Government: Research and Practice (DGOV), 1(2): 15:1–15:20, April 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3360001.

Hastings:2022:PDM

[HH22] Justine S. Hastings and Mark Howison. Predicting divertible medicaid emergency department costs. Digital Government: Research and Practice (DGOV), 3(3):19:1–19:??, July 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3548692.

Hsieh:2021:PSM

[HJY^{+}21] Hsun-Ping Hsieh, Jiawei Jiang, Tzu-Hsin Yang, Renfen Hu, and Cheng-Lin Wu. Predicting the success of mediation requests using case properties and textual information for reducing the burden on the Court. Digital Government: Research and Practice (DGOV), 2(4):30:1–30:18, October 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469233.

Honary:2020:GCB

[HMW^{+}20] Mahsa Honary, Veronica Martinez, Theodore Wlazlowski, Sumi Helal, Hans-Henning Von Oertzen, and Souroush Honary. Getting the country back to work, safely: a digital solution. Digital Government: Research and Practice (DGOV), 1(4):29:1–29:5, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3416090.

Hagen:2020:SMU

[HNSK20] Loni Hagen, Stephen Neely, Ryan Scharf, and Thomas E. Keller. Social media use for crisis and emergency risk communications during the Zika health crisis. Digital Government: Research and Practice (DGOV), 1(2):13:1–13:21, April 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3372021.

Harrison:2021:DPP

[HP21] Teresa M. Harrison and Theresa A. Pardo. Data, politics and public health: COVID-19 data-driven decision making in public discourse. Digital Government: Research and Practice (DGOV), 2(1):11:1–11:8, January 2021. CODEN ????

ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428123.

Huynh:2021:ARR

[HTH^{+}21] Trung Dong Huynh, Niko Tsakalakis, Ayah Helal, Sophie Stalla-Bourdillon, and Luc Moreau. Addressing regulatory requirements on explanations for automated decisions with provenance — a case study. Digital Government: Research and Practice (DGOV), 2(2):16e:1–16e:14, March 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436897.

Jarvis:2020:SRH

[Jar20] Jeff Jarvis. A society relearning how to talk with itself. Digital Government: Research and Practice (DGOV), 1(1): 4:1–4:10, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3352461.

Joshi:2020:SRF

[JS20] Karuna Pande Joshi and Srishty Saha. A semantically rich framework for knowledge representation of code of federal regulations. Digital Government: Research and Practice (DGOV), 1(3):21:1–21:17, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3425192.

Jagadish:2021:CBD

[JSH21] H. V. Jagadish, Julia Stoyanovich, and Bill Howe. COVID-19 brings data equity challenges to the fore. Digital Government: Research and Practice (DGOV), 2(2): 24:1–24:7, March 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3440889.

Ju:2020:TUM

[JYAM^{+}20] Wendy Ju, Sharon Yavo-Ayalon, Ilan Mandel, Federico Saldarini, Natalie Friedman, Srinath Sibi, J. D. Zamfirescu-Pereira, and Jorge Ortiz. Tracking urban mobility and occupancy under social distancing policy. Digital Government: Research and Practice (DGOV), 1(4):32:1–32:12, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3417991.

Kristensen:2023:CSL

[KA23] Kenneth Kristensen and Kim Normann Andersen. C-suite leadership of digital government. Digital Government: Research and Practice (DGOV), 4(1): 3:1–3:??, March 2023. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3580000.

Kamensky:2022:BRS

[Kam22] John M. Kamensky. Book review: Solving Public Problems — A Practical Guide to Fix Our Government and Change Our World, by Beth Simone Noveck (New Haven, CT: Yale University Press, 2021, 352 pgs.). Digital Government: Research and Practice (DGOV), 3(3):22:1–22:??, July 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542951.

Kumar:2020:SCI

[KGKS20] Munish Kumar, Surbhi Gupta, Krishan Kumar, and Monika Sachdeva. Spreading of COVID-19 in India, Italy, Japan, Spain, UK, US: a prediction using ARIMA and LSTM model. Digital Government: Research and Practice (DGOV), 1(4):24:1–24:9, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3411760.

Kalampokis:2023:UUE

[KKTT23] Evangelos Kalampokis, Nikos Karacapilidis, Dimitris Tsakalidis, and Konstantinos Tarabanis. Understanding the use of emerging technologies in the public sector: a review of Horizon 2020 projects. Digital Government: Research and Practice (DGOV), 4(1):4:1–4:??, March 2023. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580603.

Kielty:2022:CCD

[KLN22] Abigail Kielty, Amy Lee, and Michael A. Neblo. Connecting to Congress during COVID-19: Political representation and two-way crisis communication. Digital Government: Research and Practice (DGOV), 3(2): 12:1–12:??, April 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3524062.

Kleiman:2020:GMI

[KMJ20] Fernando Kleiman, Sebastiaan Meijer, and Marijn Janssen. Gaming for meaningful interactions in teleworking lessons learned during the COVID-19 pandemic from integrating gaming in virtual meetings. Digital Government: Research and Practice (DGOV), 1(4): 31:1–31:5, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3416308.

Kim:2021:BHD

[KPR21] Junhan Kim, Sun Young Park, and Lionel P. Robert. Bridging the health disparity of African Americans through conversational agents. Digital Government: Research and Practice (DGOV), 2(1):4:1–4:7, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428122.

Kumari:2021:GPR

[KS21] Moothedath Chandran Kumari and Biradar Sagar. Global pandemic and rapid new product development of medical products. Digital Government: Research and Practice (DGOV), 2(2):17:1–17:38, March 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428492.

Krieg:2021:PPU

[KSM^{+}21] Steven J. Krieg, Jennifer J.

Schnur, Jermaine D. Marshall, Matthew M. Schoenbauer, and Nitesh V. Chawla. Pandemic pulse: Unraveling and modeling social signals during the COVID-19 pandemic. Digital Government: Research and Practice (DGOV), 2(2): 19:1–19:9, March 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3431805.

Kumar:2022:CEP

[Kum22] Sachin Kumar. A critical and ethical perspective on governance and policies during COVID-19 pandemic in India. Digital Government: Research and Practice (DGOV), 3(1):

2:1–2:10, January 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3501032.

Lee:2023:HCA

[Lee23] Gwanhoo Lee. How can the artificial intelligence of things create public value? Lessons learned from use cases. Digital Government: Research and Practice (DGOV), 4(1): 5:1–5:??, March 2023. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3580604.

Lee:2021:OXB

[LL21] Roy Ka-Wei Lee and Zhi Li. Online xenophobic behavior amid the COVID-19 pandemic: a commentary. Digital Government: Research and Practice (DGOV), 2(1): 2:1–2:5, January 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428091.

Li:2022:IDL

[LL22] Yunrong Li and Gang Li. The impacts of digital literacy on citizen civic engagementevidence from China. Digital Government: Research and Practice (DGOV), 3(4):24:1– 24:??, October 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3532785.

Liu:2022:ISI

[LM22] Zhenling Liu and Liang Ma.

Introduction of the special issue on building digital government in China: New trends of building digital government in China. Digital Government: Research and Practice (DGOV), 3(4):23:1–23:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569424.

Liu:2022:PMG

[LT22] Yisheng Liu and Anying Tang.

Prediction method of government economic situation based on big data analysis. Digital Government: Research and Practice (DGOV), 3(4):27:1– 27:??, October 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3563042.

Lungati:2022:DSI

[Lun22] Angela Oduor Lungati. Data sharing and information platforms in crisis response and preparedness: Exploring the role of open data sharing platforms and collective intelligence in COVID-19 response efforts, and preparedness for future pandemics. Digital Government: Research and Practice (DGOV), 3(2):9:1–9:??, April 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3529165.

Marti:2022:CCM

[Mar22] José Martí. Crowdsourcing crisis management and democratic legitimacy. Digital Government: Research and Practice (DGOV), 3(2):15:1–15:??, April 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3532670.

Mwainyekule:2020:PEA

[MF20] Leah H. Mwainyekule and Francis B. Frimpong. The pandemic and the economy of Africa: Conflicting strategies between Tanzania and Ghana. Digital Government: Research and Practice (DGOV), 1(4): 33:1–33:8, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3419217.

Mkude:2023:OIG

[Mku23] Catherine Mkude. Open innovation in e-government: Exploring its practices in Tanzania. Digital Government: Research and Practice (DGOV), 4(1):1:1–1:??, March 2023. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571822.

Classe:2021:PPO

[MMBM21] Tadeu Moreira De Classe, Renata Mendes De Araujo, Geraldo Bonorino Xexéo, and Sean Wolfgand Matsui Siqueira. Public processes are open for play. Digital Government: Research and Practice (DGOV), 2(4):32:1–32:18, October 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474879.

Marti:2022:ISI

[MN22] José Luis Martí and Beth Simone Noveck. Introduction to the special issue on crowdlaw and emergency collective intelligence. Digital Government: Research and Practice (DGOV), 3(2):7:1–7:??, April 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550083.

Mueller:2021:UDP

[MP21] Klaus Mueller and Eric Papenhausen. Using demographic pattern analysis to predict COVID-19 fatalities on the US county level. Digital Government: Research and Practice (DGOV), 2(1):12:1–12:11, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3430196.

Mulgan:2022:GBH

[Mul22] Geoff Mulgan. Government as a brain: How can governments better understand, think, create, and remember, and avoid the traps of collective stupidity both in emergencies and normal times. Digital Government: Research and Practice (DGOV), 3(2):8:1– 8:??, April 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3528381.

Noveck:2020:DDP

[NC20] Beth Simone Noveck and Vinton G. Cerf. Digital democracy: Past, present, future an interview with Vinton G. Cerf, Vice President and Chief Internet Evangelist, Google. Digital Government: Research and Practice (DGOV), 1(1): 2:1–2:10, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3382738.

Perez:2020:CRD

[Per20] Oren Perez. Collaborative erulemaking, democratic bots, and the future of digital democracy. Digital Government: Research and Practice (DGOV), 1(1):8:1–8:13, February 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3352463.

Pitoski:2021:NAI

[PLP21] Dino Pitoski, Thomas J. Lampoltshammer, and Peter Parycek. Network analysis of internal migration in Austria. Digital Government: Research and Practice (DGOV), 2(3): 25:1–25:24, July 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3447539.

Porwol:2022:FVC

[PMM^{+}22] Lukasz Porwol, Sara S. Metcalf, J. Bradley Morrison, Soon Ae Chun, and Luis Felipe Luna-Reyes. Facilitating virtual conferences: Reflections and lessons learned in two global communities. Digital Government: Research and Practice (DGOV), 3(1): 1:1–1:12, January 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3494676.

Peng:2022:IEM

[PXW22] Jiangui Peng, Jianzheng Xu, and Yeni Wang. Innovative evaluation method of government physical education training quality based on fuzzy mathematics. Digital Government: Research and Practice (DGOV), 3(4):28:1–28:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568168.

Pekkola:2022:CMC

[PYM22] Samuli Pekkola, Maija Ylinen, and Nicholas Mavengere. Consortium of municipalities co-tailoring a governmental eservice platform: What could go wrong? Digital Government: Research and Practice (DGOV), 3(1):6:1–6:16, January 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511889.

Roy:2021:USC

[RAM^{+}21] Dibyendu Roy, Nitin Agarwal, Amitava Mukherjee, Jitendra Angne, Anees Farhan, Deven Sorathiya, R. Devarajan, and Manukumar Velayudhan Nair. Utilizing smart city cyber-physical infrastructure for tracking and monitoring pandemics like COVID-19 with the ICCC as the nerve centre. Digital Government: Research and Practice (DGOV), 2(1):8:1–8:8, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428124.

Roddiger:2021:RAD

[RBDB21] Tobias Röddiger, Michael Beigl, Dominik Dörner, and Matthias Budde. Responsible, automated data gathering for timely citizen dashboard provision during a global pandemic (COVID-19). Digital Government: Research and Practice (DGOV), 2(1):9:1–9:9, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428471.

Reynante:2021:FOC

[RDM21] Brandon Reynante, Steven P.

Dow, and Narges Mahyar. A framework for open civic design: Integrating public participation, crowdsourcing, and design thinking. Digital Government: Research and Practice (DGOV), 2(4):31:1–31:22, October 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487607.

Redmiles:2021:UCT

[Red21] Elissa M. Redmiles. User concerns 8 tradeoffs in technologyfacilitated COVID-19 response. Digital Government: Research and Practice (DGOV), 2(1):6:1–6:12, January 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3428093.

Ruprechter:2020:DHA

[RKG20] Thorsten Ruprechter, Foaad Khosmood, and Christian Guetl. Deconstructing humanassisted video transcription and annotation for legislative proceedings. Digital Government: Research and Practice (DGOV), 1(3):19:1–19:24, December 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3395316.

Sandoval-Almazan:2020:ISI

[SAUK20] Rodrigo Sandoval-Almazan, Jolien Ubacht, and Andrea Kavanaugh. Introduction to the special issue on government and social media. Digital Government: Research and Practice (DGOV), 1(2): 10:1–10:3, April 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381940.

Sandoval-Almazan:2020:SAF

[SAVC20] Rodrigo Sandoval-Almazan and David Valle-Cruz. Sentiment analysis of Facebook users reacting to political campaign posts. Digital Government: Research and Practice (DGOV), 1(2):12:1–12:13, April 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3382735.

Scholl:2020:DGL

[Sch20a] Hans J. Scholl. Digital government: Looking back and ahead on a fascinating domain of research and practice. Digital Government: Research and Practice (DGOV), 1(1): 7:1–7:12, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3352682.

Schuler:2020:CTS

[Sch20b] Douglas Schuler. Can technology support democracy? Digital Government: Research and Practice (DGOV), 1(1):3:1– 3:14, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3352462.

Schuler:2022:COA

[Sch22] Douglas Schuler. Computing as oppression: Authoritarian technology poses a worldwide threat. Digital Government: Research and Practice (DGOV), 3(4):29:1–29:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568400.

Sclove:2020:DTI

[Scl20] Richard Sclove. Democracy and technology: an interview with Richard Sclove from Beth Simone Noveck. Digital Government: Research and Practice (DGOV), 1(1): 5:1–5:6, February 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3368273.

Stachofsky:2022:CVS

[SGC22] Justin Stachofsky, Assefaw Gebremedhin, and Robert E. Crossler. Cast to vote: a socio-technical network analysis of an election smartphone application. Digital Government: Research and Practice (DGOV), 3(1):3:1–3:17, January 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3501031.

Sgouros:2022:PIN

[Sgo22] Nikitas Marinos Sgouros. Politika: Implementing a novel meta-simulation methodology for public policy design on the Web. Digital Government: Research and Practice (DGOV), 3(4):30:1–30:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3568167.

Sierra:2022:CVC

[Sie22] Jose M. Martinez Sierra. Constitutions vs. COVID-19: CrowdLaw as an alternative in crisis and recovery. Digital Government: Research and Practice (DGOV), 3(2): 14:1–14:??, April 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524063.

Steinfeld:2020:MFD

[SLO20] Nili Steinfeld and Azi Lev-On. MPs on Facebook: Differences between members of coalition and opposition. Digital Government: Research and Practice (DGOV), 1(2): 11:1–11:14, April 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3363415.

Spatola:2021:WRC

[SM21] Nicolas Spatola and Karl F.

Macdorman. Why real citizens would turn to artificial leaders. Digital Government: Research and Practice (DGOV), 2(3): 26:1–26:24, July 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3447954.

Smyth:2020:LRD

[Smy20] Barry Smyth. Lockdowns & rebounds: a data analysis of what happens next. Digital Government: Research and Practice (DGOV), 1(4):27:1– 27:7, December 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3411762.

Smyth:2021:EFB

[Smy21]  Barry Smyth.  Estimat-

ing the fatality burden of SARS-CoV2. Digital Government: Research and Practice (DGOV), 2(2):21:1–21:8, March 2021. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436997.

Suran:2022:BGS

[SPK^{+}22] Shweta Suran, Vishwajeet Pattanaik, Ralf Kurvers, Carina Antonia Hallin, Anna De Liddo, Robert Krimmer, and Dirk Draheim. Building global societies on collective intelligence: Challenges and opportunities. Digital Government: Research and Practice (DGOV), 3(4):31:1–31:??, October 2022. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3568169.

Tremblay-Cantin:2023:GSA

[TCMCAK23] Claudie-Ann Tremblay-Cantin, Sehl Mellouli, Mustapha Cheikh-Ammar, and Hager Khechine. E-government service adoption by citizens: a literature review and a high-level model of influential factors. Digital Government: Research and Practice (DGOV), 4(1):2:1–2:??, March 2023. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580369.

Tian:2022:SUS

[TH22] Zhen Tian and Huaichen Hu. Studies of urban safety improvement for anti-corrosion smokestack protection based on LOP strategies. Digital Government: Research and Practice (DGOV), 3(4):25:1– 25:??, October 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3528622.

Tam:2021:ICS

[TRH21] Tracy Tam, Asha Rao, and Joanne Hall. The invisible COVID-19 small business risks: Dealing with the cyber-security aftermath. Digital Government: Research and Practice (DGOV), 2(2): 23:1–23:8, March 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3436807.

Valtolina:2022:LGW

[VF22] Stefano Valtolina and Daniele Fratus. Local government Websites accessibility: Evaluation and finding from Italy. Digital Government: Research and Practice (DGOV), 3(3): 17:1–17:??, July 2022. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/10.1145/3528380.

Wu:2020:EPA

[WH20] Qunfang Wu and Yun Huang.

Examining police agencies’ dialogic accounting practices in Facebook conversations. Digital Government: Research and Practice (DGOV), 1(2): 16:1–16:17, April 2020. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372022.

Yaqub:2021:TDC

[Yaq21] Ussama Yaqub. Tweeting during the Covid-19 pandemic:

Sentiment analysis of Twitter messages by President Trump. Digital Government: Research and Practice (DGOV), 2(1): 1:1–1:7, January 2021. CO-DEN ???? ISSN 2691-199X (print), 2639-0175 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428090.

Yaqub:2020:LBS

[YSP^{+}20] Ussama Yaqub, Nitesh Sharma, Rachit Pabreja, Soon Ae Chun, Vijayalakshmi Atluri, and Jaideep Vaidya. Locationbased sentiment analyses and visualization of Twitter election data. Digital Government: Research and Practice (DGOV), 1(2):14:1–14:19, April 2020. CODEN ???? ISSN 2691-199X (print), 2639-0175 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3339909.